# On Space-Filling Input Design for Nonlinear Dynamic Model Learning: A Gaussian Process Approach

### v1

@@ V-B Nonlinear mass-spring-damper system example
Regarding the parametric input signal, a periodic multisine excitation according to (3) is selected with $N_{f}=92$ excited frequencies between $[1,10]$ Hz, a sampling frequency $f_{s}=100$ Hz, and $N=1024$ points per period for this example and with $k_{\mathrm{min}}=11$ and $k_{\mathrm{max}}=102$. The optimization is performed by tuning both the amplitudes and the phases of each frequency component, denoted as $\theta=\mathrm{vec}(A_{l},\phi_{l}),l\in\mathbb{I}_{0}^{N_{f}-1}$. Specifically, the initialization of the phase $\phi_{l}$ is realized by sampling from a uniform distribution $\mathcal{U}(0,2\pi)$, and the initial value for the amplitude $A_{l}$ is selected as 100 $N$. In addition, considering the output limitations of the actuators, we set the upper bound of $A_{l}$ as 200 $N$. To simplify the optimization problem, we assume that all frequency components share the same amplitude. The region of interest is selected as $\tilde{\mathcal{Z}}:=[-400,400]\times[-2,2]\times[-20,20]$, with $M=512$ uniformly distributed anchor points $\tilde{z}_{i}$ with $i\in\mathbb{I}_{i}^{M}$. The kernel hyperparameters are chosen as $\sigma_{\mathrm{f}}^{2}=10$ and $\Lambda^{1/2}=\mathrm{diag}(120,0.6,6)$.
A 3D depiction of the considered domain can be observed in Fig. 3. Although the initial data is highly concentrated around the center of the domain, after optimization a clear space-filling behavior can be observed. It is important to emphasize that these results were observed using a signal with $N=1024$ samples while exciting only a limited range of frequencies. Table I provides a quantitative evaluation of the value of $\rho(\mathcal{D}_{N}(\theta))$ and $\rho(\mathcal{D}_{\mathrm{I}})$ (mean value over 10 Monte Carlo simulations) under different sizes of $\mathcal{D}_{\mathrm{I}}$ before and after optimization with the normalization weight $Q=\mathrm{diag}\left(\frac{1}{400^{2}},\frac{1}{2^{2}},\frac{1}{20^{2}}\right)$. It can be observed that minimizing the proposed GP-based space-filling promoting cost function (14) likewise leads to the minimization of the classical criterion (15). Even though the optimization is nonlinear in this scenario, which prevents the optimization cost from reaching 0 and results in $\rho(\mathcal{D}_{N}(\theta))$ being slightly larger than $\rho(\mathcal{D}_{\mathrm{I}})$ in the $M=512$ case, our proposed method still achieves adequate space-filling behavior.
@@
Num. of anchor points $\rho(\mathcal{D}_{N}(\theta))$ $\rho(\mathcal{D}_{\mathrm{I}})$
$M=512$ Init. 1.2756 0.2449
@@ VI Conclusion
This letter presents an effective identification-oriented GP-based space-filling input design strategy for nonlinear dynamical model learning. A GP prior is placed on the latent joint input-state space, resulting in an innovative space-filling promotion cost function from a Bayesian perspective. Theoretical analysis has been provided to demonstrate the equivalence between the convergence of the cost function and the expected space-filling behavior. The introduced approach provides new perspectives into the input design for data-driven model learning, especially in nonlinear cases lacking systematic design methods and with the flexible parameterization of the input signal. The effectiveness of the proposed strategy has been analyzed and demonstrated by numerical simulations.

### v2

@@ V-B Nonlinear mass-spring-damper system example
This example illustrates the effectiveness of the proposed experiment design approach for nonlinear systems, even when Condition 4 is violated by limiting the frequency and amplitude spectrum range of the input signal.
@@
A multisine excitation (3) is considered with $N_{f}\!=\!92$ excited frequencies between $[1,10]$ Hz, a sampling frequency $f_{s}\!=\!100$ Hz, $N\!=\!1024$ points per period, and with $k_{\mathrm{min}}\!=\!11$ and $k_{\mathrm{max}}\!=\!102$. Both the amplitudes and the phases of each frequency component are optimized, denoted as $\theta\!=\!\mathrm{vec}(A_{l},\phi_{l}),l\in\mathbb{I}_{0}^{N_{f}-1}$. The phase $\phi_{l}$ is initialized by sampling from a uniform distribution $\mathcal{U}(0,2\pi)$, and the initial amplitude value $A_{l}$ is selected as 100 N. In addition, to limit the input signal power, we set the upper bound of $A_{l}$ to 200 N. To simplify the optimization problem, we assume that all frequency components share the same amplitude. The region of interest is selected as $\tilde{\mathcal{Z}}\!:=\![-400,400]\!\times\![-2,2]\!\times\![-20,20]$, with $M\!=\!512$ uniformly distributed anchor points $\tilde{z}_{i}$ with $i\in\mathbb{I}_{1}^{M}$. The kernel hyperparameters are chosen as $\sigma_{\mathrm{f}}^{2}\!=\!10$ and $\Lambda^{1/2}\!=\!\mathrm{diag}(120,0.6,6)$.
A 3D depiction of the considered domain is shown in Fig. 3. Although the initial data is highly concentrated around the center of the domain, after optimization, a clear space-filling behavior can be observed. It is important to emphasize that these results were observed using a $N\!=\!1024$ while exciting only a limited frequency range. Table I provides a quantitative evaluation of $\rho(\mathcal{D}_{N}(\theta))$ and $\rho(\mathcal{D}_{\mathrm{I}})$ (mean value over 10 Monte Carlo simulations) under different sizes of $\mathcal{D}_{\mathrm{I}}$ before and after optimization with the normalization weight $Q\!=\!\mathrm{diag}\left(\frac{1}{400^{2}},\frac{1}{2^{2}},\frac{1}{20^{2}}\right)$. It can be observed that minimizing the proposed GP-based space-filling promoting cost function (13) leads to the minimization of the classical criterion (14). As the input is restricted to a user-defined frequency range and amplitude spectrum, we expect local minima in the proposed cost due to the nonlinear system dynamics, a nonzero optimal cost can be expected. While $\rho(\mathcal{D}_{N}(\theta))$ is slightly larger than $\rho(\mathcal{D}_{\mathrm{I}})$ in the $M\!=\!512$ case, the method still significantly improves the space-filling property of the dataset over the unoptimized input, validating its robustness even beyond the strict theoretical conditions. We further compare the proposed method with a Schroeder phase multisine, a common choice in experimental design. The corresponding $\rho(\mathcal{D}_{N}^{\text{MS}})\!=\!1.1811$, which is lower than the unoptimized input but significantly higher than the values obtained by the proposed approach. This demonstrates the improvement achieved by the proposed strategy over conventional input design choices.
@@
Num. of anchor points $\rho(\mathcal{D}_{N}(\theta))$ $\rho(\mathcal{D}_{\mathrm{I}})$ $\rho(\mathcal{D}_{N}^{\text{MS}})$
$M=512$ Init. 1.2756 0.2449 1.1811
@@ VI Conclusion
This paper proposes an effective GP-based space-filling input design strategy for nonlinear system identification. A GP prior is placed on the latent joint input-state space, resulting in a novel space-filling promoting cost function from a Bayesian perspective under flexible parameterization of the input signal. Theoretical analysis demonstrates the equivalence between the convergence of the cost function and the expected space-filling behavior. Effectiveness of the proposed strategy has been analyzed and demonstrated on experiment design examples, showing the space-covering nature of the obtained data.